# Multiplicities of weakly graded families of ideals

### v2

@@ 2. $(ii)$
$(i)$ First we consider $\overline{R[\mathcal{I}]}=\overline{R[\mathcal{J}]}$. Let $\overline{R[\mathcal{I}]}=R[\mathcal{C}]$ where $\mathcal{C}=\{C_{n}\}$ is a divisorial filtration. Note that $I_{n}\subset J_{n}\subset C_{n}$ for all $n\geq 1$. By Lemma 2.9, we have $R[\mathcal{C}]=\overline{R[\mathcal{C}]}$. Hence
  1. $(a)$
    $(I_{n}:K)\subset(J_{n}:K)\subset(C_{n}:K)$ for all $n\geq 1$, and
  2. $(b)$
    by [9, Theorem 5.1], we have $e(\mathcal{J})=e(\mathcal{I})=e(\mathcal{C})$.
Using $(i)$ and $(iii)$ of Proposition 4.1, we have
$e(\mathcal{C})=\lim\limits_{n\to\infty}d!\ell_{R}(R/(C_{n}:K))/n^{d}$$\leq$ $\lim\limits_{n\to\infty}d!\ell_{R}(R/(J_{n}:K))/n^{d}$
$\leq$ $\lim\limits_{n\to\infty}d!\ell_{R}(R/(I_{n}:K))/n^{d}\leq e(\mathcal{I})=e(\mathcal{C}).$
@@ 2. $(b)$
Now we prove the converse. Let $\lim\limits_{n\to\infty}\ell_{R}(R/(I_{n}:K))/n^{d}=\lim\limits_{n\to\infty}\ell_{R}(R/(J_{n}:K))/n^{d}$. Suppose $\overline{R[\mathcal{I}]}\neq\overline{R[\mathcal{J}]}$. Let $\overline{R[\mathcal{I}]}=R[\mathcal{C}]$ where $\mathcal{C}=\{C_{n}\}$ is a divisorial filtration. By Lemma 2.9, we have $R[\mathcal{C}]=\overline{R[\mathcal{C}]}$. Then by [9, Theorem 5.1], we have $e(\mathcal{I})=e(\mathcal{C})$.
@@
Since $\overline{R[\mathcal{I}]}\neq\overline{R[\mathcal{J}]}$, by [7, Theorem 14.4], we have $e(\mathcal{I})\neq e(\mathcal{J})$. Now $I_{n}\subset J_{n}$ for all $n\geq 1$ imply
$e(\mathcal{J})<e(\mathcal{I})=e(\mathcal{C}).$ (4)
Since $I_{n}\subset C_{n}$ for all $n\geq 1$, using $(i)$ and $(iii)$ of Proposition 4.1 and equation (4), we get
$\lim\limits_{n\to\infty}d!\ell_{R}(R/(J_{n}:K))/n^{d}$$\leq$ $e(\mathcal{J})$
$<$ $e(\mathcal{C})=\lim\limits_{n\to\infty}d!\ell_{R}(R/(C_{n}:K))/n^{d}$
$\leq$ $\lim\limits_{n\to\infty}d!\ell_{R}(R/(I_{n}:K))/n^{d}$
which contradicts the assumption.
@@
$(ii)$ Let $\overline{R[\mathcal{I}]}=R[\mathcal{C}]$ and $\overline{R[\mathcal{J}]}=R[\mathcal{T}]$ where $\mathcal{C}=\{C_{n}\}$ and $\mathcal{T}=\{T_{n}\}$ are integral divisorial filtrations. Note that $I_{n}\subset C_{n}$ and $J_{n}\subset T_{n}$ for all $n\geq 1$. Using Lemma 2.9, we have $\overline{R[\mathcal{I}]}=R[\mathcal{C}]=\overline{R[\mathcal{C}]}$ and $\overline{R[\mathcal{J}]}=R[\mathcal{T}]=\overline{R[\mathcal{T}]}$. Thus by [9, Theorem 5.1], we have $e(\mathcal{I})=e(\mathcal{C})$ and $e(\mathcal{J})=e(\mathcal{T})$. Therefore using part $(i)$ and part $(iii)$ of Proposition 4.1, we have
$e(\mathcal{I})=e(\mathcal{C})=\lim\limits_{n\to\infty}d!\ell_{R}(R/(C_{n}:K))/n^{d}\leq\lim\limits_{n\to\infty}d!\ell_{R}(R/(I_{n}:K))/n^{d}\leq e(\mathcal{I})$
and
$e(\mathcal{J})=e(\mathcal{T})=\lim\limits_{n\to\infty}d!\ell_{R}(R/(T_{n}:K))/n^{d}\leq\lim\limits_{n\to\infty}d!\ell_{R}(R/(J_{n}:K))/n^{d}\leq e(\mathcal{J}).$
Consider the filtration =$\mathcal{H}=\{I_{n}J_{n}\}$. If we show that
$\lim\limits_{n\to\infty}d!\ell_{R}(R/(I_{n}:K)(J_{n}:K))/n^{d}=e(\mathcal{H})$
then the result follows from [7, Theorem 14.5].
@@
By part $(iii)$ of Proposition 4.1, there exist positive integers $w$ and $w^{\prime}$ such that $(I_{wn}:K)\subset I_{w(n-1)}$ and $(J_{w^{\prime}n}:K)\subset J_{w^{\prime}(n-1)}$ for all $n\geq 1$. Let $u=ww^{\prime}$. Then
$I_{un}\subset(I_{un}:K)=(I_{ww^{\prime}n}:K)\subset I_{w(w^{\prime}n-1)}\subset I_{ww^{\prime}(n-1)}=I_{u(n-1)}$
and
$J_{un}\subset(J_{un}:K)=(J_{ww^{\prime}n}:K)\subset J_{w^{\prime}(wn-1)}\subset J_{ww^{\prime}(n-1)}=J_{u(n-1)}$
for all $n\geq 1$. Therefore
$e(\mathcal{H})$ $=$ $\lim\limits_{n\to\infty}d!\ell_{R}(R/I_{u(n-1)}J_{u(n-1)})/{(un)}^{d}$
$\leq$ $\lim\limits_{n\to\infty}d!\ell_{R}(R/(I_{un}:K)(J_{un}:K))/{(un)}^{d}$
$\leq$ $\lim\limits_{n\to\infty}d!\ell_{R}(R/(I_{un}J_{un})/{(un)}^{d}=e(\mathcal{H}).$
@@ 1. $(i)$
  2. $(iii)$
    Let =$\mathcal{I}=\{I_{n}\}$ be a Noetherian filtration of ideals in $R$ with $\operatorname{grade}(I_{1})\geq 1$ and $K$ be an $\mathfrak{m}$-primary ideal in $R$ with $I_{1}\subset K$. Then
@@ 2. $(iii)$
    In particular, if $I$ is an ideal in $R$ with $\operatorname{grade}(I)\geq 1$ and $K$ is an $\mathfrak{m}$-primary ideal in $R$ with $I\subset K$ then
@@
$(ii)$ Let =$\mathcal{I}=\{I_{n}\}$ be a Noetherian filtration of ideals in $R$ with $\operatorname{grade}(I_{1})\geq 1$ and $K$ be an $\mathfrak{m}$-primary ideal in $R$ with $I_{1}\subset K$. Then by [2, Proposition 2.4] and [28, Theorem 3.4], there exists $r\in{\mathbb{Z}}_{>0}$ such that $\mathcal{I}$ satisfies $A(r)$ condition. Hence using [5, Theorem 6.1], we have $\varepsilon(\mathcal{I})$ exists and by $(i)$, $\lim\limits_{n\to\infty}d!\ell_{R}(H_{\mathfrak{m}}^{0}(R/(I_{n}:K)))/{n}^{d}$ exists.
Since $\mathcal{I}$ is a Noetherian filtration, By [1, Proposition 3, page 159], there exists an integer $e\geq 1$ such that $I_{en}=I_{e}^{n}$ for all $n\geq 1$. Therefore using $(i)$, we have
@@
Since $R$ is an analytically unramified local ring, by [24], $\{\overline{I^{n}}\}$ is a Noetherian filtration. By [5, Corollary 6.3], $\varepsilon(I)=\lim\limits_{n\to\infty}d!\ell_{R}({(I^{n})}^{\operatorname{sat}}/{I^{n}})/{n}^{d}$ exists and by [16, Proposition 2.1], $\varepsilon(\mathcal{I})=\lim\limits_{n\to\infty}d!\ell_{R}(\overline{I^{n}}^{\operatorname{sat}}/\overline{I^{n}})/{n}^{d}$ exists and $\varepsilon(\mathcal{I})=\varepsilon(I)$. Thus we get the required result. ∎
@@
By [3, Lemma 4.2], we have $\lim\limits_{n\to\infty}d!\ell_{R}(({I_{e}^{n}}:\mathfrak{m})/I_{e}^{n})/{n}^{d}=0$. Hence by replacing $K$ by $\mathfrak{m}$ in equations (6) and (4), for any Noetherian filtration =$\mathcal{I}=\{I_{n}\}$ (without the assumption that $\operatorname{grade}(I_{1})\geq 1$) in an analytically unramified local ring $(R,\mathfrak{m})$, we get $\lim\limits_{n\to\infty}d!\ell_{R}(H_{\mathfrak{m}}^{0}(R/({I_{n}}:\mathfrak{m})))/{n}^{d}=\varepsilon(\mathcal{I})$.

### v3

@@ 2. $(ii)$
Let $X=\lim\limits_{n\to\infty}d!\ell_{R}(R/(I_{n}:K)(J_{n}:K))/n^{d}$, $Y=\lim\limits_{n\to\infty}d!\ell_{R}(R/(I_{n}:K)/n^{d}$ and $Z=\lim\limits_{n\to\infty}d!\ell_{R}(R/(J_{n}:K))/n^{d}$.
@@
$(i)$ Let $\overline{R[\mathcal{I}]}=R[\mathcal{C}]$ where $\mathcal{C}=\{C_{n}\}$ is a divisorial filtration. By Lemma 2.9, we have $R[\mathcal{C}]=\overline{R[\mathcal{C}]}$ and by [9, Theorem 5.1], we have $e(\mathcal{I})=e(\mathcal{C})$.
Suppose $\overline{R[\mathcal{I}]}=\overline{R[\mathcal{J}]}$. Then $I_{n}\subset J_{n}\subset C_{n}$ for all $n\geq 1$ and using $(i)$ and $(iii)$ of Proposition 4.2, we get,
$e(\mathcal{C})=\lim\limits_{n\to\infty}d!\ell_{R}(R/(C_{n}:K))/n^{d}\leq Z\leq Y\leq e(\mathcal{I})=e(\mathcal{C})$
which implies $Y=Z$.
@@
Now we prove the converse. Let $Y=Z$. Suppose $\overline{R[\mathcal{I}]}\neq\overline{R[\mathcal{J}]}$. By [7, Theorem 14.4], we have $e(\mathcal{I})\neq e(\mathcal{J})$. Now $I_{n}\subset J_{n}$ for all $n\geq 1$, imply $e(\mathcal{J})<e(\mathcal{I})=e(\mathcal{C})$. Since $I_{n}\subset C_{n}$ for all $n\geq 1$, using $(i)$ and $(iii)$ of Proposition 4.2, we get
$Z\leq e(\mathcal{J})<e(\mathcal{C})=\lim\limits_{n\to\infty}d!\ell_{R}(R/(C_{n}:K))/n^{d}\leq Y$
which contradicts the assumption $Y=Z$.
@@
$(ii)$ Due to part $(iii)$ of Theorem 3.2, it is enough to prove that $X^{1/d}\geq Y^{1/d}+Z^{1/d}$. Let $\overline{R[\mathcal{I}]}=R[\mathcal{C}]$ and $\overline{R[\mathcal{J}]}=R[\mathcal{T}]$ where $\mathcal{C}=\{C_{n}\}$ and $\mathcal{T}=\{T_{n}\}$ are integral divisorial filtrations. Note that $I_{n}\subset C_{n}$ and $J_{n}\subset T_{n}$ for all $n\geq 1$. By Lemma 2.9, we have $\overline{R[\mathcal{I}]}=R[\mathcal{C}]=\overline{R[\mathcal{C}]}$ and $\overline{R[\mathcal{J}]}=R[\mathcal{T}]=\overline{R[\mathcal{T}]}$. Therefore using [9, Theorem 5.1], part $(i)$ and part $(iii)$ of Proposition 4.2, we have
$e(\mathcal{I})=e(\mathcal{C})=\lim\limits_{n\to\infty}d!\ell_{R}(R/(C_{n}:K))/n^{d}\leq Y\leq e(\mathcal{I})\mbox{ and}$
$e(\mathcal{J})=e(\mathcal{T})=\lim\limits_{n\to\infty}d!\ell_{R}(R/(T_{n}:K))/n^{d}\leq Z\leq e(\mathcal{J}).$
Hence $Y=e(\mathcal{C})$ and $Z=e(\mathcal{T}).$ Therefore by [7, Theorem 14.5], it is enough to show that $X\geq e(\mathcal{H})$ where =$\mathcal{H}=\{C_{n}T_{n}\}$. Note that
$X=\lim\limits_{n\to\infty}d!\ell_{R}(R/(I_{n}:K)(J_{n}:K))/n^{d}\geq\lim\limits_{n\to\infty}d!\ell_{R}(R/(C_{n}:K)(T_{n}:K))/n^{d}.$
@@
Thus it is enough to show that $\lim\limits_{n\to\infty}d!\ell_{R}(R/(C_{n}:K)(T_{n}:K))/n^{d}\geq e(\mathcal{H})$. By part $(iii)$ of Proposition 4.2, there exist positive integers $w$ and $w^{\prime}$ such that $(C_{wn}:K)\subset C_{w(n-1)}$ and $(T_{w^{\prime}n}:K)\subset T_{w^{\prime}(n-1)}$ for all $n\geq 1$. Let $u=ww^{\prime}$. Then for all $n\geq 1$,
$C_{un}\subset(C_{un}:K)=(C_{ww^{\prime}n}:K)\subset C_{w(w^{\prime}n-1)}\subset C_{ww^{\prime}(n-1)}=C_{u(n-1)}\mbox{ and}$
$T_{un}\subset(T_{un}:K)=(T_{ww^{\prime}n}:K)\subset T_{w^{\prime}(wn-1)}\subset T_{ww^{\prime}(n-1)}=T_{u(n-1)}.$
Therefore
$e(\mathcal{H})$ $=$ $\lim\limits_{n\to\infty}d!\ell_{R}(R/C_{u(n-1)}T_{u(n-1)})/{(u(n-1))}^{d}$
$\leq$ $\lim\limits_{n\to\infty}\big{[}\Big{(}d!\ell_{R}(R/(C_{un}:K)(T_{un}:K))/{(un)}^{d}\Big{)}\Big{(}{(un)}^{d}/{(u(n-1))}^{d}\Big{)}\big{]}$
$=$ $\lim\limits_{n\to\infty}d!\ell_{R}(R/(C_{n}:K)(T_{n}:K))/n^{d}.$
@@ 1. $(i)$
  2. $(ii)$
    Let =$\mathcal{I}=\{I_{n}\}$ be a Noetherian filtration of ideals in $R$ with $\operatorname{grade}(I_{1})\geq 1$ and $K$ be an $\mathfrak{m}$-primary ideal in $R$. Then
@@ 2. $(ii)$
    In particular, if $I$ is an ideal in $R$ with $\operatorname{grade}(I)\geq 1$ and $K$ is an $\mathfrak{m}$-primary ideal in $R$ then
@@
$(ii)$ Let =$\mathcal{I}=\{I_{n}\}$ be a Noetherian filtration of ideals in $R$ with $\operatorname{grade}(I_{1})\geq 1$. If $I_{n}=R$ for all $n\geq 1$, then we get the required result. Suppose $I_{1}\neq R$. By [2, Proposition 2.4] and [28, Theorem 3.4], there exists $r\in{\mathbb{Z}}_{>0}$ such that $\mathcal{I}$ satisfies $A(r)$ condition. Hence using [5, Theorem 6.1], we have $\varepsilon(\mathcal{I})$ exists and by $(i)$, $\lim\limits_{n\to\infty}d!\ell_{R}(H_{\mathfrak{m}}^{0}(R/(I_{n}:K)))/{n}^{d}$ exists. By Remark 4.1, there exists an integer $e\geq 1$ such that $I_{e}\subset K$ and $I_{en}=I_{e}^{n}$ for all $n\geq 1$. Therefore using $(i)$, we have
@@
Since $R$ is an analytically unramified local ring, by [24], =$\mathcal{I}=\{\overline{I^{n}}\}$ is a Noetherian filtration. By [5, Corollary 6.3], $\varepsilon(I)=\lim\limits_{n\to\infty}d!\ell_{R}({(I^{n})}^{\operatorname{sat}}/{I^{n}})/{n}^{d}$ exists and by [16, Proposition 2.1], $\varepsilon(\mathcal{I})=\lim\limits_{n\to\infty}d!\ell_{R}(\overline{I^{n}}^{\operatorname{sat}}/\overline{I^{n}})/{n}^{d}$ exists and $\varepsilon(\mathcal{I})=\varepsilon(I)$. Thus we get the required result. ∎
@@
By [3, Lemma 4.2], we have $\lim\limits_{n\to\infty}d!\ell_{R}(({I_{e}^{n}}:\mathfrak{m})/I_{e}^{n})/{n}^{d}=0$. Hence by replacing $K$ by $\mathfrak{m}$ in part $(ii)$ of Theorem 4.4, for any Noetherian filtration =$\mathcal{I}=\{I_{n}\}$ (without the assumption that $\operatorname{grade}(I_{1})\geq 1$) in an analytically unramified local ring $(R,\mathfrak{m})$, we get $\lim\limits_{n\to\infty}d!\ell_{R}(H_{\mathfrak{m}}^{0}(R/({I_{n}}:\mathfrak{m})))/{n}^{d}=\varepsilon(\mathcal{I})$.